# The Enhanced Physics-Informed Kolmogorov–Arnold Networks: Applications of Newton’s Laws in Financial Deep Reinforcement Learning (RL) Algorithms

### v1

@@ Abstract
Deep Reinforcement Learning (DRL), a subset of machine learning focused on sequential decision-making, has emerged as a powerful approach for tackling financial trading problems. In finance, DRL is commonly used either to generate discrete trade signals or to determine continuous portfolio allocations. In this work, we propose a novel reinforcement learning framework for portfolio optimization that incorporates Physics-Informed Kolmogorov-Arnold Networks (PIKANs) into several DRL algorithms. The approach replaces conventional multilayer perceptrons with Kolmogorov-Arnold Networks (KANs) in both actor and critic components—utilizing learnable B-spline univariate functions to achieve parameter-efficient and more interpretable function approximation. During actor updates, we introduce a physics-informed regularization loss that promotes second-order temporal consistency between observed return dynamics and the action-induced portfolio adjustments. The proposed framework is evaluated across three equity markets—China, Vietnam, and the United States, covering both emerging and developed economies. Across all three markets, PIKAN-based agents consistently deliver higher cumulative and annualized returns, superior Sharpe and Calmar ratios, and more favorable drawdown characteristics compared to both standard DRL baselines and classical online portfolio-selection methods. This yields more stable training, higher Sharpe ratios, and superior performance compared to traditional DRL counterparts. The approach is particularly valuable in highly dynamic and noisy financial markets, where conventional DRL often suffers from instability and poor generalization.
@@ 5 Model Design
Mathematical Formulations in the General Framework
@@
      Actor EMA: $\mu_{\text{actor}}\leftarrow\beta\mu_{\text{actor}}+(1-\beta)\cdot\text{mean}(Q(s_{t},a_{t}))$, $\sigma^{2}_{\text{actor}}\leftarrow\beta\sigma^{2}_{\text{actor}}+(1-\beta)\cdot\text{var}(Q(s_{t},a_{t}))$.
@@ 6.2 Performance metrics
To evaluate the effectiveness of the proposed portfolio allocation strategies, we adopt a multi-dimensional set of quantitative indicators. These metrics assess performance across total profitability, risk-adjusted returns, and downside protection, following established benchmarks in financial deep reinforcement learning Jiang & Liang [2016]. The primary measures—**Cumulative Return Rate (CRR)**, **Annualized Return (AR)**, **Annualized Volatility (AV)**, **Sharpe Ratio (SR)**, **Maximum Drawdown (MDD)**, and **Calmar Ratio (CR)**—are defined below.
@@ 8 Conclusion
CRediT authorship contribution statement
First Author: Conceptualization, Methodology, Software, Writing – original draft, Writing – review & editing. Second Author: Methodology, Software, Writing – review & editing. Last Author: Supervision, Methodology, Validation.
Declaration of Competing Interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Acknowledgements
Collate acknowledgements in a separate section at the end of the article before the references and do not, therefore, include them on the title page, as a footnote to the title or otherwise. List here those individuals who provided help during the research (e.g., providing language help, writing assistance or proof reading the article, etc.).
Data availability
Provide a link to your data if available.

### v2

# The Enhanced Physics-Informed Kolmogorov–Arnold Networks: Applications of Newton’s Laws in Financial Deep Reinforcement Learning Algorithms

@@ Abstract
Under the field of quantitative finance, Deep Reinforcement Learning (DRL) is commonly used either to generate discrete trade signals or to determine continuous portfolio allocations. In this work, we propose a novel reinforcement learning framework for portfolio optimization that incorporates Physics-Informed Kolmogorov-Arnold Networks (PIKANs) into several DRL algorithms. The approach replaces conventional multilayer perceptrons with Kolmogorov-Arnold Networks (KANs) in both actor and critic components—utilizing learnable B-spline univariate functions to achieve parameter-efficient and more interpretable function approximation. During actor updates, we introduce a physics-informed regularization loss that promotes second-order temporal consistency between observed return dynamics and the action-induced portfolio adjustments. The proposed framework is evaluated across three equity markets—the United States, Vietnam, and China, covering both emerging and developed economies. Across all three markets, PIKAN-based agents consistently deliver higher cumulative and annualized returns, superior Sharpe and Calmar ratios, and more favorable drawdown characteristics compared to both standard DRL baselines and classical online portfolio-selection methods. This yields more stable training, higher Sharpe ratios, and superior performance compared to traditional DRL counterparts. The approach is particularly valuable in highly dynamic and noisy financial markets, where conventional DRL often suffers from instability and poor generalization.
@@ 5 Model Design
Mathematical Formulations in the General Framework:
@@
      Actor EMA:
      $\mu_{\text{actor}}$ $=\beta\mu_{\text{actor}}+(1-\beta)\cdot\text{mean}(Q(s_{t},a_{t}))$
      $\sigma^{2}_{\text{actor}}$ $=\beta\sigma^{2}_{\text{actor}}+(1-\beta)\cdot\text{var}(Q(s_{t},a_{t}))$
@@ 6.2 Performance metrics
To evaluate the effectiveness of the proposed portfolio allocation strategies, we adopt a multi-dimensional set of quantitative indicators. These metrics assess performance across total profitability, risk-adjusted returns, and downside protection, following established benchmarks in financial deep reinforcement learning Jiang & Liang [2016]. The primary measures—Cumulative Return Rate (CRR), Annualized Return (AR), Annualized Volatility (AV), Sharpe Ratio (SR), Maximum Drawdown (MDD), and Calmar Ratio (CR)—are defined below.